HIGH-RESOLUTION SPECTROSCOPY AND ANALYSIS OF THE ν_{3}/2ν_{4} DYAD OF CF_{4}

\underline{V. BOUDON}, Institut Carnot de Bourgogne, UMR 5209 CNRS–Université de Bourgogne, 9 Av. A. Savary, BP 47870, F-21078 Dijon Cedex, France; A. DOMANSKAYA, Institute of Physics, St. Petersburg University, 198504 St. Petersburg, Russia; C. MAUL, Institut für Physikalische und Theoretische Chemie der Technischen Universität Braunschweig, D-38106 Braunschweig, Germany; R. GEORGES, Institut de Physique de Rennes, UMR 6251 CNRS–Universit de Rennes 1, 263 Av. Général Leclerc, F-35042 Rennes Cedex, France; J. MITCHELL, University of Arkansas, 226 Physics Building, Fayetteville AR 72701, USA.

CF_{4} is a strong greenhouse gas of both anthropogenic and natural origin^{a}. However, high-resolution infrared spectroscopy of this molecule has received only a limited interest up to now. The public databases only contain cross-sections for this species, but no detailed line list. We reinvestigate here the strongly absorbing ν_{3} region around 7.3 μm. Two new Fourier transform infrared spectra at a 0.003 cm^{−1} resolution have been recorded: i) a room-temperature spectrum in a static cell with a 5 mb pressure and ii) a supersonic expansion jet spectrum at a 15 K estimated temperature. Following the work of Gabard et al.^{b}, we perform a simultaneous analysis of both the ν_{3} and 2ν_{4} bands since a strong Coriolis interaction occurs between them, perturbing the ν_{3} R-branch rotational clusters around J = 20. As in Ref. b, we also include ν_{3} − ν_{3} microwave data in the fit. The analysis is performed thanks to the XTDS and SPVIEW programs^{c}. Compared to Ref. b, the present work extends the analysis up to higher J values (56 instead of 32). Absorption intensities are estimated thanks to the dipole moment derivative value of D. Papoušek et al.^{d} and compare well with the experiment. The rotational energy surfaces for the ν_{3}/2ν_{4} dyad are also examined in order to understand the distribution of rovibrational levels.

aD. R. Worton, W. T. Sturges, L. K. Gohar et al., Environ. Sci. Technol. 41, 2184–2189 (2007).

bT. Gabard, G. Pierre and M. Takami, Mol. Phys. 85, 735–744 (1995).

cCh. Wenger, V. Boudon, M. Rotger, M. Sanzharov and J.-P. Champion, J. Mol. Spectrosc., 251 102–113 (2008).

dD. Papoušek, Z. Papoušková and D. P. Chong, J. Phys. Chem. 99, 15387–15395 (1995).